OpenStudy (anonymous):

The path of a soccer ball is modeled by the function h(x) = -0.005x2 + 0.25x, where h is the height in meters and x is the horizontal distance that the ball travels in meters. What is the maximum height that the ball reaches? ___________ m.

OpenStudy (raden):

h will be max when x=-b/2a, (a=-0.005, b=0.25) so, x=-0.25/(2*-0.005) = 25 now, find the value of h(25)